3. The following table lists the nominal federal spending on nondefense-related education and training (in million dollars) and CPI (price index) in 1962, 1965, 200, and 2007. 1962 19652004 2007 Nominal Spending$1,306 1,959 $85,798 90,701 CPI30.2 31.5 188.9 207.3 (a) Compare the change in nominal spending from 1962 to 1965 and that from 2001 to 2004, which increase was larger? (b) Compare the change in real spending from 1962 to 1965 and that from 2004 to 2007, which increase was larger? Note that you need to calculate the real spending first.
Answer #1

(a) Change in nominal spending

1062-65: (1,959 / 1,306) - 1 = 1.5 - 1 = 0.5 = 50%

2004-07: (90,701 / 85,798) - 1 = 1.0571 - 1 = 0.0571 = 5.71%

Increase in nominal spending was higher between 1962-65.

(b)

Real spending = (Nominal spending / CPI) x 100

1962: ($1,306 / 30.2) x 100 = $4,325

1965: ($1,959 / 31.5) x 100 = $6,219

2004: ($85,798 / 188.9) x 100 = $45,420

2007: ($90,701 / 207.3) x 100 = $43,753

Increase in real spending, 1962-65: (6,219 / 4,325) - 1 = 1.4379 - 1 = 0.4379 = 43.79%

Increase in real spending, 2004-07: (43,753 / 45,420) - 1 = 0.9633 - 1 = - 0.0367 = - 3.67%

Increase in real spending was higher between 1962-75.